# Efficient Learning for Entropy-regularized Markov Decision Processes via Multilevel Monte Carlo

### v1

@@ 3 Numerical Experiments
  • The MLMCu estimator, with appropriately chosen hyperparameters, achieves polynomial complexity, whereas the MLMCb estimator exhibits quasi-polynomial complexity.
@@ 3.2 Numerical Results
According to Theorem 2.6, we expect a power law relationship between these two quantities for MLMCu. This is confirmed by the straight lines in Figure 2(b). In contrast, it seems that the MLMCb estimator does not exhibit a power law, as can be clearly seen for $K=2$ in Figure 2. This shows that the MLMCb error indeed suffers from a quasi-polynomial complexity, confirming Theorem 2.3. The behaviour stems from the intrinsic bias of the plain Monte Carlo.
@@ 4 Analysis of the Simple Iterative MC Estimator
$\sigma_{\mathbf{T}}(s,a)\coloneqq\mathrm{Var}(T^{0}Q_{0}(S^{0}_{s,a}))^{\frac{1}{2}},\quad\delta_{\mathbf{T}}(s,a)\coloneqq\left|\mathbb{E}\left[T^{0}Q^{\star}(S^{0}_{s,a})-TQ^{\star}(S^{0}_{s,a})\right]\right|.$
Then $\sigma_{\mathbf{T}_{K}}(s,a)\leq L^{\prime}K^{-1/2}$ and $\delta_{\mathbf{T}_{K}}(s,a)\leq L^{\prime}K^{-1/2}$, where $L^{\prime}=\tau\left[e^{(\beta-\alpha)/\tau}-1\right]$.
@@
$\exp\left[-\tau^{-1}\hat{T}_{K}^{0}Q_{0}(S^{0}_{s,a})\right],\exp\left[-\tau^{-1}\hat{T}_{K}^{0}Q^{\star}(S^{0}_{s,a})\right],\exp\left[-\tau^{-1}TQ^{\star}(S^{0}_{s,a})\right]\in\left[e^{-\beta/\tau},e^{-\alpha/\tau}\right].$
The function $g:x\mapsto-\tau\log x$ is Lipschitz on $\left[e^{-\beta/\tau},e^{-\alpha/\tau}\right]$, with a Lipschitz constant given by $\tau e^{\beta/\tau}$. Therefore, we have
$\sigma^{2}_{\mathbf{T}_{K}}(s,a)$ $=\mathrm{Var}\left[g\left(\frac{1}{K}\sum_{k=1}^{K}e^{-Q(S^{0}_{s,a},A^{(0,k)})/\tau}\right)\right]$
$\leq\mathbb{E}\left|g\left(\frac{1}{K}\sum_{k=1}^{K}e^{-Q(S^{0}_{s,a},A^{(0,k)})/\tau}\right)-g\left(\mathbb{E}e^{-Q(S^{0}_{s,a},A^{(0,1)})/\tau}\right)\right|^{2}$
$\leq\frac{\tau^{2}e^{2\beta/\tau}}{K}\mathrm{Var}\left[\exp(-Q(S_{s,a}^{0},A^{(0,1)})/\tau)\right]$
$\leq\frac{\tau^{2}e^{2\beta/\tau}}{K}\left[e^{-\alpha/\tau}-e^{-\beta/\tau}\right]^{2}=\frac{(L^{\prime})^{2}}{K}.$
Similarly, we have
$\left\|\exp\left[-\hat{T}_{K}^{0}Q^{\star}(S^{0}_{s,a})/\tau\right]-\exp\left[-TQ^{\star}(S^{0}_{s,a})/\tau\right]\right\|^{2}_{L^{2}}$
$=\frac{1}{K}\mathrm{Var}\left[\exp\left(-Q^{*}(S^{0}_{s,a},A^{0})/\tau\right)-\exp\left(-TQ^{\star}(S^{0}_{s,a})/\tau\right)\right].$
This implies
$\delta^{2}_{\mathbf{T}_{K}}(s,a)$ $\leq\left(\mathbb{E}\left|\hat{T}^{0}_{K}Q^{\star}(S^{0}_{s,a})-TQ^{\star}(S^{0}_{s,a})\right|\right)^{2}$
$\leq\left\|\hat{T}_{K}^{0}Q^{\star}(S^{0}_{s,a})/\tau-TQ^{\star}(S^{0}_{s,a})\right\|^{2}_{L^{2}}$
$\leq\tau^{2}e^{2\beta/\tau}\left\|\exp\left[-\hat{T}_{K}^{0}Q^{\star}(S^{0}_{s,a})/\tau\right]-\exp\left[-TQ^{\star}(S^{0}_{s,a})/\tau\right]\right\|^{2}_{L^{2}}$
$=\frac{\tau^{2}e^{2\beta/\tau}}{K}\mathrm{Var}\left[\exp\left(-Q^{*}(S^{0}_{s,a},A^{0})/\tau\right)-\exp\left(-TQ^{\star}(S^{0}_{s,a})/\tau\right)\right]$
$\leq\frac{\tau^{2}e^{2\beta/\tau}}{K}\left[e^{-\alpha/\tau}-e^{-\beta/\tau}\right]^{2}=\frac{(L^{\prime})^{2}}{K}.$
@@
$\|Q_{n}^{\theta}(s,a)-Q^{\star}(s,a)\|_{L^{2}}\leq\sqrt{\mathrm{Var}Q_{n}^{\theta}(s,a)}+\left|\mathbb{E}\left[Q^{\theta}_{n}(s,a)-Q^{\star}(s,a)\right]\right|.$
@@
$\leq\gamma L\sup_{s,a}\|Q^{\theta}_{n-1}(s,a)-Q^{\star}(s,a)\|_{L^{2}}+\gamma\|\delta_{\mathbf{T}}\|_{\infty},$
@@
By Lemma 4.4, we know that $\|\delta_{\mathbf{T}}\|_{\infty}\leq K^{-1/2}L^{\prime}$, where
$L^{\prime}=\tau\exp\left(\tau^{-1}\beta\right)\left[\exp\left(\beta\right)-\exp\left(\alpha\right)\right]$
is a constant depending only on $c_{\min},c_{\max},\gamma,\tau$, hence the final bound is
$E_{n}\leq\frac{\gamma\sqrt{C}}{\sqrt{M}(1-\gamma L)}+\frac{\gamma L^{\prime}}{\sqrt{K}(1-\gamma L)}+(\gamma L)^{n}E_{0},$
@@
$n=\left\lceil\frac{\log\varepsilon-\log 3E_{0}}{\log\gamma L}\right\rceil,M=\left\lceil 9\frac{\gamma^{2}C}{(1-\gamma L)^{2}\varepsilon^{2}}\right\rceil,K=\left\lceil 9\frac{\gamma^{2}(L^{\prime})^{2}}{(1-\gamma L)^{2}\varepsilon^{2}}\right\rceil.$
We get $E_{n}\leq\varepsilon/3+\varepsilon/3+\varepsilon/3=\varepsilon$ and the sample complexity is of order $\varepsilon^{\frac{-4\log\varepsilon}{\log\gamma L}(1+o(1))}$. ∎
@@ 5.2 Distributional Properties of the General MLMC Estimator
We state a lemma which ensures that $\theta$ only contributes as an index in the definition of the general MLMC estimator given by (2.7) and (2.8) and state some measurability properties of the general MLMC estimator in the framework of admissible stochastic operators.
@@
    $(Q^{\theta^{\prime}}_{n,M,\mathbf{T}})_{\theta^{\prime}\in\Theta}$ are identically distributed.
@@ 5.3 Analysis of the Monte Carlo Error
$\mathrm{Var}(Q_{n,M,\mathbf{T}}^{\theta}(s,a))=\mathrm{Var}(Q_{n,M,\mathbf{T}}^{0}(s,a))\leq\mathrm{Var}(\hat{Q}_{n,M,\mathbf{T}}^{0}(s,a)),$
@@
$\mathrm{Var}(\hat{Q}_{n,M,\mathbf{T}}^{0}(s,a))\leq\frac{\gamma^{2}\sigma_{\mathbf{T}}(s,a)^{2}}{M^{n}}+\sum_{l=1}^{n}\frac{(\gamma L)^{2}}{M^{n-l}}\left\|Q_{l,M,\mathbf{T}}^{(0,l)}(S_{s,a}^{0},A^{0})-Q_{l-1,M,\mathbf{T}}^{(0,-l)}(S_{s,a}^{0},A^{0})\right\|^{2}_{L^{2}},$
where $\sigma_{\mathbf{T}}(s,a)^{2}=\mathrm{Var}(T^{0}Q_{0}(S^{0}_{s,a}))$.
@@
$\mathrm{Var}(Q^{0}_{n,M,\mathbf{T}}(s,a))$ $\leq\mathrm{Var}(\hat{Q}^{0}_{n,M,\mathbf{T}}(s,a))$
$=\frac{\gamma^{2}}{M^{n}}\mathrm{Var}[T^{(0,0,1)}Q_{0}(S^{0}_{s,a})]$
$+\sum_{l=1}^{n-1}\frac{\gamma^{2}}{M^{n-l}}\mathrm{Var}\left[T^{(0,l,1)}Q_{l,M,\mathbf{T}}^{(0,l,1)}(S_{s,a}^{(0,l,1)})-T^{(0,l,1)}Q_{l,M,\mathbf{T}}^{(0,-l,1)}(S_{s,a}^{(0,l,1)})\right].$
@@
$\mathrm{Var}(\hat{Q}_{n,M,\mathbf{T}}^{0}(s,a))$ $=\frac{\gamma^{2}}{M^{n}}\sigma_{\mathbf{T}}(s,a)^{2}$
$+\sum_{l=1}^{n}\frac{\gamma^{2}}{M^{n-l}}\|T^{(0,l,1)}Q_{l,M,\mathbf{T}}^{(0,l,1)}(S^{(0,l,1)}_{s,a})-T^{(0,l,1)}Q_{l-1,M,\mathbf{T}}^{(0,-l,1)}(S^{(0,l,1)}_{s,a})\|_{L^{2}}^{2}$
$\leq\frac{\gamma^{2}}{M^{n}}\sigma_{\mathbf{T}}(s,a)^{2}$
$+\sum_{l=1}^{n}\frac{(\gamma L)^{2}}{M^{n-l}}\left\|Q_{l,M,\mathbf{T}}^{(0,l,1)}(S_{s,a}^{(0,l,1)},A^{(0,l,1,1)})-Q_{l-1,M,\mathbf{T}}^{(0,-l,1)}(S_{s,a}^{(0,l,1)},A^{(0,l,1,1)})\right\|^{2}_{L^{2}}.$
@@ 5.4 Analysis of the Truncation Error
$\delta^{\mathrm{trunc}}_{n,M,\mathbf{T}}(s,a)^{2}\leq\max\left\{\mathbb{P}(\hat{Q}^{\theta}_{n,M,\mathbf{T}}(s,a)<\alpha),\mathbb{P}(\hat{Q}^{\theta}_{n,M,\mathbf{T}}(s,a)>\beta)\right\}\mathrm{Var}(\hat{Q}^{\theta}_{n,M,\mathbf{T}}(s,a)).$
@@
$\delta^{\mathrm{trunc}}_{n}(s,a)=\left|\mathbb{E}\left[\mathbf{1}_{\hat{Q}_{n}(s,a)<\alpha}(\alpha-\hat{Q}_{n}(s,a))+\mathbf{1}_{\hat{Q}_{n}(s,a)>\beta}(\beta-\hat{Q}_{n}(s,a))\right]\right|.$
@@
$\mathbb{E}\left[1_{\hat{Q_{n}}(s,a)<\alpha}(\alpha-\hat{Q}_{n}(s,a))\right]^{2}$ $\leq\mathbb{E}\left[1_{\hat{Q_{n}}(s,a)<\alpha}(\mathbb{E}[\hat{Q}_{n}(s,a)]-\hat{Q}_{n}(s,a))\right]^{2}$
$\leq\mathbb{P}(\hat{Q}_{n}(s,a)<\alpha)\mathrm{Var}(\hat{Q}_{n}(s,a)).$
@@
$\mathbb{E}\left[1_{\hat{Q}_{n}(s,a)>\beta}(\beta-\hat{Q}_{n}(s,a))\right]^{2}\leq\mathbb{P}(\hat{Q}_{n}(s,a)>\beta)\mathrm{Var}(\hat{Q}_{n}(s,a)).$
@@
$\delta^{\mathrm{trunc}}_{n}(s,a)^{2}\leq\max\left\{\mathbb{P}(\hat{Q}_{n}(s,a)<\alpha),\mathbb{P}(\hat{Q}_{n}(s,a)>\beta)\right\}\mathrm{Var}(\hat{Q}_{n}(s,a)).$

### v3

@@ 3 Numerical Experiments
  • The MLMCu estimator, with appropriately chosen hyperparameters, achieves polynomial complexity, whereas the MLMCb estimator exhibits super-polynomial complexity.
@@ 3.2 Numerical Results
According to Theorem 2.6, we expect a power law relationship between these two quantities for MLMCu. This is confirmed by the straight lines in Figure 2(b). In contrast, it seems that the MLMCb estimator does not exhibit a power law, as can be clearly seen for $K=2$ in Figure 2. This shows that the MLMCb error indeed suffers from a super-polynomial complexity, confirming Theorem 2.3. This behaviour stems from the intrinsic bias of the plain Monte Carlo average for approximating $T$ (1.4).
@@ 4 Analysis of the Simple Iterative MC Estimator
$\sigma_{\mathbf{T}}(s,a)\coloneqq\mathrm{Var}\left(T^{0}Q_{0}\left(S^{0}_{s,a}\right)\right)^{\frac{1}{2}},\quad\delta_{\mathbf{T}}(s,a)\coloneqq\left|\mathbb{E}\left[T^{0}Q^{\star}\left(S^{0}_{s,a}\right)-TQ^{\star}\left(S^{0}_{s,a}\right)\right]\right|.$
Then $\sigma_{\mathbf{T}_{K}}(s,a)\leq L^{\prime}K^{-1/2}$ and $\delta_{\mathbf{T}_{K}}(s,a)\leq(L^{\prime})^{2}(2\tau K)^{-1}$, where $L^{\prime}=\tau\left(e^{(\beta-\alpha)/\tau}-1\right)$.
@@
$\exp\left(-\hat{T}_{K}^{0}Q_{0}(S^{0}_{s,a})/\tau\right),\exp\left(-\hat{T}_{K}^{0}Q^{\star}(S^{0}_{s,a})/\tau\right),\exp\left(-TQ^{\star}(S^{0}_{s,a})/\tau\right)\in\left[e^{-\beta/\tau},e^{-\alpha/\tau}\right].$
The function $g:x\mapsto-\tau\log x$ is Lipschitz on $\left[e^{-\beta/\tau},e^{-\alpha/\tau}\right]$, with a Lipschitz constant given by $\tau e^{\beta/\tau}$. Let
$m(S_{s,a}^{0})=\int_{\mathcal{A}}\exp(-Q^{\star}(S_{s,a}^{0},b)/\tau)\mu({\mathrm{d}}b),\quad\Sigma_{K}=\frac{1}{K}\sum_{k=1}^{K}\exp(-Q^{\star}(S_{s,a}^{0},A^{(0,k)})/\tau).$
We have
$\sigma^{2}_{\mathbf{T}_{K}}(s,a)$ $=\mathrm{Var}\left[g\left(\Sigma_{K}\right)\right]$
$\leq\mathbb{E}\left|g\left(\Sigma_{K}\right)-g\left(\mathbb{E}\exp\left(-Q(S^{0}_{s,a},A^{(0,1)})/\tau\right)\right)\right|^{2}$
$\leq\frac{\tau^{2}e^{2\beta/\tau}}{K}\mathrm{Var}\left[\exp\left(-Q\left(S_{s,a}^{0},A^{(0,1)}\right)/\tau\right)\right]$
$\leq\frac{\tau^{2}e^{2\beta/\tau}}{K}\left(e^{-\alpha/\tau}-e^{-\beta/\tau}\right)^{2}=\frac{(L^{\prime})^{2}}{K}.$
Since $g$ is smooth on $(0,\infty)$ and $m(S_{s,a}^{0}),\Sigma_{K}\in[e^{-\beta/\tau},e^{-\alpha/\tau}]$, we perform a Taylor expansion with Lagrange remainder
$g(\Sigma_{K})=g(m(S_{s,a}^{0}))+g^{\prime}(m(S^{0}_{s,a}))(\Sigma_{K}-m(S^{0}_{s,a}))+\frac{g^{\prime\prime}(\xi_{K})}{2}(\Sigma_{K}-m(S^{0}_{s,a}))^{2},$
where $\xi_{K}\in[\Sigma_{K},m(S^{0}_{s,a})]$. Moreover, since $g^{\prime\prime}$ is strictly monotone and continuous, $\xi_{K}$ is defined uniquely and continuously as a function of $\Sigma_{K},m(S^{0}_{s,a})$. In particular $\xi_{K}$ is a random variable. Hence
$\delta_{\mathbf{T}_{K}}(s,a)=\left|\mathbb{E}\left[g^{\prime}(m(S^{0}_{s,a}))(\Sigma_{K}-m(S^{0}_{s,a}))+\frac{g^{\prime\prime}(\xi_{K})}{2}(\Sigma_{K}-m(S^{0}_{s,a}))^{2}\right]\right|.$
The first order term can be rewritten using the tower property
$\mathbb{E}\left[g^{\prime}(m(S^{0}_{s,a}))(\Sigma_{K}-m(S^{0}_{s,a}))\right]=\mathbb{E}\left[g^{\prime}(m(S_{s,a}^{0}))\mathbb{E}\left[(\Sigma_{K}-m(S^{0}_{s,a}))\mid S_{s,a}^{0}\right]\right],$
and by independence of $A^{(0,k)}$ and $S^{0}_{s,a}$, we have $\mathbb{E}\left[(\Sigma_{K}-m(S^{0}_{s,a}))\mid S_{s,a}^{0}\right]=0$. Hence
$\delta_{\mathbf{T}_{K}}(s,a)$ $=\left|\mathbb{E}\left[\frac{g^{\prime\prime}(\xi_{K})}{2}(\Sigma_{K}-m(S^{0}_{s,a}))^{2}\right]\right|$
$\leq\sup_{x\in[e^{-\beta/\tau},e^{-\alpha/\tau}]}\frac{\tau}{2x^{2}}\mathbb{E}\left[\left(\Sigma_{K}-m(S^{0}_{s,a})\right)^{2}\right]$
$=\frac{\tau e^{2\beta/\tau}}{2K}\mathbb{E}\left[\left(\exp(-Q^{\star}(S^{0}_{s,a},A^{(0,1)}))-m(S_{s,a}^{0})\right)^{2}\right]$
$\leq\frac{\tau e^{2\beta/\tau}}{2K}\left(e^{-\alpha/\tau}-e^{-\beta/\tau}\right)^{2}=\frac{(L^{\prime})^{2}}{2\tau K}.$
@@
$\left\|Q_{n}^{\theta}(s,a)-Q^{\star}(s,a)\right\|_{L^{2}}\leq\sqrt{\mathrm{Var}Q_{n}^{\theta}(s,a)}+\left|\mathbb{E}\left[Q^{\theta}_{n}(s,a)-Q^{\star}(s,a)\right]\right|.$
@@
$\leq\gamma L\sup_{s,a}\left\|Q^{\theta}_{n-1}(s,a)-Q^{\star}(s,a)\|_{L^{2}}+\gamma\right\|\delta_{\mathbf{T}}\|_{\infty},$
@@
By Lemma 4.4, we know that $\|\delta_{\mathbf{T}}\|_{\infty}\leq(L^{\prime})^{2}(2\tau K)^{-1}$, where $L^{\prime}=\tau(L-1)$ is a constant depending only on $c_{\min},c_{\max},\gamma,\tau$, hence the final bound is
$E_{n}\leq\frac{\gamma\sqrt{C}}{\sqrt{M}(1-\gamma L)}+\frac{\gamma(L^{\prime})^{2}}{2\tau K(1-\gamma L)}+(\gamma L)^{n}E_{0},$
@@
$n=\left\lceil\frac{\log\varepsilon-\log 3E_{0}}{\log\gamma L}\right\rceil,M=\left\lceil 9\frac{\gamma^{2}C}{(1-\gamma L)^{2}\varepsilon^{2}}\right\rceil,K=\left\lceil 3\frac{\gamma(L^{\prime})^{2}}{2\tau(1-\gamma L)\varepsilon}\right\rceil.$
We get $E_{n}\leq\varepsilon/3+\varepsilon/3+\varepsilon/3=\varepsilon$ and the sample complexity is of order $\varepsilon^{\frac{-3\log\varepsilon}{\log\gamma L}(1+o(1))}$. ∎
@@ 5.2 Distributional Properties of the General MLMC Estimator
We state a lemma which ensures that $\theta$ only contributes as an index in the definition of the general MLMC estimator given by (2.7) and state some measurability properties of the general MLMC estimator in the framework of admissible stochastic operators.
@@
    $\left(Q^{\theta^{\prime}}_{n,M,\mathbf{T}}\right)_{\theta^{\prime}\in\Theta}$ are identically distributed.
@@ 5.3 Analysis of the Monte Carlo Error
$\mathrm{Var}\left(Q_{n,M,\mathbf{T}}^{\theta}(s,a)\right)=\mathrm{Var}\left(Q_{n,M,\mathbf{T}}^{0}(s,a)\right)\leq\mathrm{Var}\left(\hat{Q}_{n,M,\mathbf{T}}^{0}(s,a)\right),$
@@
$\mathrm{Var}\left(\hat{Q}_{n,M,\mathbf{T}}^{0}(s,a)\right)\leq\frac{\gamma^{2}\sigma_{\mathbf{T}}(s,a)^{2}}{M^{n}}+\sum_{l=1}^{n}\frac{(\gamma L)^{2}}{M^{n-l}}\left\|Q_{l,M,\mathbf{T}}^{(0,l)}\left(S_{s,a}^{0},A^{0}\right)-Q_{l-1,M,\mathbf{T}}^{(0,-l)}\left(S_{s,a}^{0},A^{0}\right)\right\|^{2}_{L^{2}},$
where $\sigma_{\mathbf{T}}(s,a)^{2}=\mathrm{Var}\left(T^{0}Q_{0}(S^{0}_{s,a})\right)$.
@@
$\mathrm{Var}\left(Q^{0}_{n,M,\mathbf{T}}(s,a)\right)$ $\leq\mathrm{Var}\left(\hat{Q}^{0}_{n,M,\mathbf{T}}(s,a)\right)$
$=\frac{\gamma^{2}}{M^{n}}\mathrm{Var}\left[T^{(0,0,1)}Q_{0}\left(S^{0}_{s,a}\right)\right]$
$+\sum_{l=1}^{n-1}\frac{\gamma^{2}}{M^{n-l}}\mathrm{Var}\left[T^{(0,l,1)}Q_{l,M,\mathbf{T}}^{(0,l,1)}\left(S_{s,a}^{(0,l,1)}\right)-T^{(0,l,1)}Q_{l,M,\mathbf{T}}^{(0,-l,1)}\left(S_{s,a}^{(0,l,1)}\right)\right].$
@@
$\mathrm{Var}\left(\hat{Q}_{n,M,\mathbf{T}}^{0}(s,a)\right)$
$\leq\frac{\gamma^{2}}{M^{n}}\sigma_{\mathbf{T}}(s,a)^{2}+\sum_{l=1}^{n}\frac{\gamma^{2}}{M^{n-l}}\left\|T^{(0,l,1)}Q_{l,M,\mathbf{T}}^{(0,l,1)}\left(S^{(0,l,1)}_{s,a}\right)-T^{(0,l,1)}Q_{l-1,M,\mathbf{T}}^{(0,-l,1)}\left(S^{(0,l,1)}_{s,a}\right)\right\|_{L^{2}}^{2}$
$\leq\frac{\gamma^{2}}{M^{n}}\sigma_{\mathbf{T}}(s,a)^{2}+\sum_{l=1}^{n}\frac{(\gamma L)^{2}}{M^{n-l}}\left\|Q_{l,M,\mathbf{T}}^{(0,l,1)}\left(S_{s,a}^{(0,l,1)},A^{(0,l,1,1)}\right)-Q_{l-1,M,\mathbf{T}}^{(0,-l,1)}\left(S_{s,a}^{(0,l,1)},A^{(0,l,1,1)}\right)\right\|^{2}_{L^{2}}.$
@@ 5.4 Analysis of the Truncation Error
$\delta^{\mathrm{trunc}}_{n,M,\mathbf{T}}(s,a)^{2}\leq\max\left\{\mathbb{P}\left(\hat{Q}^{\theta}_{n,M,\mathbf{T}}(s,a)<\alpha\right),\mathbb{P}\left(\hat{Q}^{\theta}_{n,M,\mathbf{T}}(s,a)>\beta\right)\right\}\mathrm{Var}\left(\hat{Q}^{\theta}_{n,M,\mathbf{T}}(s,a)\right).$
@@
$\delta^{\mathrm{trunc}}_{n}(s,a)=\left|\mathbb{E}\left[\mathbf{1}_{\hat{Q}_{n}(s,a)<\alpha}\left(\alpha-\hat{Q}_{n}(s,a)\right)+\mathbf{1}_{\hat{Q}_{n}(s,a)>\beta}\left(\beta-\hat{Q}_{n}(s,a)\right)\right]\right|.$
@@
$\mathbb{E}\left[1_{\hat{Q_{n}}(s,a)<\alpha}\left(\alpha-\hat{Q}_{n}(s,a)\right)\right]^{2}$ $\leq\mathbb{E}\left[1_{\hat{Q_{n}}(s,a)<\alpha}\left(\mathbb{E}\left[\hat{Q}_{n}(s,a)\right]-\hat{Q}_{n}(s,a)\right)\right]^{2}$
$\leq\mathbb{P}\left(\hat{Q}_{n}(s,a)<\alpha\right)\mathrm{Var}\left(\hat{Q}_{n}(s,a)\right).$
@@
$\mathbb{E}\left[1_{\hat{Q}_{n}(s,a)>\beta}\left(\beta-\hat{Q}_{n}(s,a)\right)\right]^{2}\leq\mathbb{P}\left(\hat{Q}_{n}(s,a)>\beta\right)\mathrm{Var}\left(\hat{Q}_{n}(s,a)\right).$
@@
$\delta^{\mathrm{trunc}}_{n}(s,a)^{2}\leq\max\left\{\mathbb{P}\left(\hat{Q}_{n}(s,a)<\alpha\right),\mathbb{P}\left(\hat{Q}_{n}(s,a)>\beta\right)\right\}\mathrm{Var}\left(\hat{Q}_{n}(s,a)\right).$